Find general equation of (D^2+2D+5)y=34sinxcos
Using method of undetermined coefficientts and d-operator methods.
1
Expert's answer
2020-05-11T10:24:47-0400

Let (D2+2D+5)y = 34 sinx cosx = 17 sin(2x) equation 1

The auxiliary equation is given by

D2+2D+5 = 0

D= -1+2i, -1-2i

The complimentary function of equation1 is e-x [c1 cos(2x) + c2 sin(2x) ] equation 2


D-OPERATOR METHOD

particular integral =  "\frac{17}{D^{2}+2D+5}" sin2x =


"\frac{17}{2D+1} sin2x"


"(2D−1) \frac{17}{4D^{2}-1}sin2x"

(1-2D) sin2x = -4cos2x + sin2x


so, the general solution of equation is

e-x [c1 cos(2x) + c2 sin(2x)] -4 cos2x + sin2x


UNDETERMINED COEFFICIENT METHOD

Let trial solution of equation1 be z= A cos(2x) + B sin(2x) equation3

Dz= -2A sin(2x) + 2B cos(2x) equation4

D2z = -4A cos(2x) - 4B sin(2x) equation5 substituting equations 3,4 and 5 in equation1

(A+4B) cos(2x) + (B-4A) sin(2x) = 17 sin(2x)

comparing like terms on both sides,

A=-4 , B=1

so, z= -4 cos(2x) + sin(2x)


so, the general solution of equation is

e-x [c1 cos(2x) + c2 sin(2x)] -4 cos2x + sin2x